What are the vector components of the following: Round off your answer to two decimal numbers.

a. G = 10m, 150°

b. L = 10m, 90°

c. A = 10m, 360°

d. D = 10m, 36° N of W


1
Expert's answer
2021-10-02T16:38:33-0400
"G_x=10\cos{150}=-8.66\ m\\G_y=10\sin{150}=5\ m\\
L_x=10\cos{90}=0\ m\\L_y=10\sin{90}=10\ m\\
A_x=10\cos{360}=10\ m\\A_y=10\sin{360}=0\ m\\
D_x=10\cos{36}=8.09\ m\\D_y=10\sin{36}=5.88\ m\\"